OpenStudy (anonymous):

you have a 3 card deck containing a king, queen, and a jack. you draw a random card then, without putting it back, you draw a second random card from the ones that are left. what is the probability that you will draw exactly one queen?

OpenStudy (anonymous):

Well, you can't draw two, now can you?

OpenStudy (anonymous):

Best way to say this is P(Q on first draw ) = 1/3 given failure + 2/3 1/2 = 1/3 total = 2/3

OpenStudy (shane_b):

Nevermind...I think I got the wording now. I think it's asking what the odds are that your second card will be a queen. The odds of the first card being a queen is 1 in 3. The odds of the second card being a queen (if I'm not mistaken) would be: \[(\frac{2}{3})(\frac{1}{2})=\frac{2}{6}=\frac{1}{3}\]

OpenStudy (anonymous):

Seems good. Notice by symmetry the odds should be equal to get any particular card on the second draw, hence P = 1/3.